### v1

Abstract
Accurate molecular property prediction is essential in drug discovery and related fields. However, existing graph neural networks (GNNs) often struggle to simultaneously capture both local and global molecular structures. In this work, we propose a Multi-Level Fusion Graph Neural Network (MLFGNN) that integrates Graph Attention Networks and a novel Graph Transformer to jointly model local and global dependencies. In addition, we incorporate molecular fingerprints as a complementary modality and introduce a mechanism of interaction between attention to adaptively fuse information across representations. Extensive experiments on multiple benchmark datasets demonstrate that MLFGNN consistently outperforms state-of-the-art methods in both classification and regression tasks. Interpretability analysis further reveals that the model effectively captures task-relevant chemical patterns, supporting the usefulness of multi-level and multi-modal fusion in molecular representation learning.
keywords:
American Chemical Society,
[…]
1 Introduction
Molecular property prediction is a fundamental task in chem-informatics and plays a crucial role in various stages of drug discovery, including virtual screening1, lead optimization2, and toxicity assessment3. Accurately predicting physicochemical and biological properties of molecules can significantly reduce the cost and time associated with experimental validation and clinical trials4. Traditionally, this task relied heavily on handcrafted features and rule-based systems, which often struggled to generalize across diverse chemical spaces5. With the growing availability of molecular data and the advancement of computational methods, there has been increasing interest in data-driven approaches, particularly those based on machine learning, to model complex structure-property relationships more effectively6.
In recent years, deep learning (DL) has achieved remarkable success in a variety of domains, such as computer vision 7 and natural language processing 8, largely driven by improvements in model architectures and the availability of large-scale data. These advances have inspired a wave of applications in scientific and biomedical fields, including atomic charge prediction 9, molecular property modeling 10, chemical reaction forecasting 11, and molecular generation 12. In particular, deep learning has shown great promise in molecular property prediction by learning generalizable representations directly from raw molecular data, thereby overcoming the limitations of handcrafted descriptors.
Among deep learning methods, graph neural networks (GNNs) have become the dominant paradigm for molecular modeling due to their ability to operate directly on molecular graph structures, where atoms are represented as nodes and bonds as edges. A variety of GNN architectures have been proposed for this task 13, 14, 15, 16. Among them, Graph Attention Networks (GATs) leverage attention mechanisms to adaptively weigh neighboring node features, making them effective for capturing local structural information 17, 18. In contrast, Graph Transformers incorporate global self-attention, enabling each node to interact with all others regardless of topological distance, which facilitates modeling global dependencies 19, 20, 21. Despite recent progress, existing models often fall short in addressing key challenges in molecular representation learning. For example, the well-known activity cliff phenomenon—where minor structural changes, such as the replacement of a single atom or functional group, can lead to dramatic shifts in biological activity—requires models that can simultaneously capture fine-grained local variations and global structural dependencies. Local-only models may overlook long-range effects, while global-only models risk missing chemically subtle but critical substructures. Moreover, feature representations such as molecular fingerprints, while chemically informative, may contain irrelevant or redundant elements that reduce prediction robustness. Compounding these challenges, different downstream tasks (e.g., solubility estimation vs. toxicity classification) often require attention to different aspects of molecular structure. These issues underscore the need for a unified framework capable of flexibly integrating diverse structural cues and adapting its focus based on task-specific demands. To address this issue, we design a multi-level fusion strategy that integrates both local and global information within the graph representation. Rather than statically combining the outputs of different modules 22, we introduce a learnable mechanism that adaptively balances their contributions based on task-specific needs.
In addition to structural information from molecular graphs, incorporating complementary modalities has shown promise for further enhancing predictive performance. Some approaches integrate molecular fingerprints 23, 24, motif-level features 25, 26, sequence representations 27, 3D conformational data 28, or even multiple sources simultaneously 29. Among these, molecular fingerprints are particularly valuable because they encode chemically meaningful substructures—such as functional groups, ring systems, or pharmacophores that are often directly linked to molecular activity. In comparison, other features may lack interpretability, contain redundant information, or introduce noise that is irrelevant to downstream tasks. Therefore, we choose molecular fingerprints as the complementary modality to be fused with graph-based representations. However, fingerprint vectors can also include irrelevant or noisy features that may harm predictive performance. Prior studies have largely overlooked this issue 23, 24. To overcome it, we introduce a cross-attention mechanism that enables the model to adaptively aggregate and filter features from both modalities, ensuring that only task-relevant information contributes to the final molecular representation.
In this study, we propose a Multi-Level Fusion Neural Network for molecular property prediction, which integrates both intra-graph (local–global) and inter-modal (graph–fingerprint) fusion. The graph branch combines local and global structural cues through the integration of GATs and a newly designed Graph Transformer. The fingerprint branch introduces domain-informed substructure features. A final cross-attention layer performs adaptive aggregation and feature selection across modalities. Extensive experiments on multiple benchmark datasets demonstrate the effectiveness, interpretability, and generalizability of our approach.
2 MATERIALS AND METHODS
2.1 Problem statement
SMILES is a textual representation of molecular structures that encodes atoms and bonds as ASCII strings. Each SMILES string $x_{i}$ can be converted into a molecular graph $G_{i}=(\mathcal{V}_{i},\mathcal{E}_{i})$, where $\mathcal{V}_{i}$ is the set of atoms and $\mathcal{E}_{i}$ is the set of chemical bonds. Given a data set $D=\{(x_{i},y_{i})|x_{i}\in X,y_{i}\in P\}$, where $X$ denotes the set of SMILES strings and $P$ represents the corresponding molecular properties, our aim is to predict molecular properties from structure. Let $M=\{G|G_{i}\in G\}$ denote the set of molecular graphs derived from the dataset. The objective is to leverage both the SMILES representations and their corresponding molecular graphs to predict the associated molecular properties. To achieve this objective, we employ deep learning techniques to train model parameters by minimizing a predefined loss function, thereby enabling more accurate prediction of molecular properties.
2.2 The architecture of MLFGNN
Algorithm 1 The Multi-Level Fusion Graph Neural Network
1: Input: SMILES string $S$, Layer $L_{1}$, Layer $L_{2}$
2: Output: Final prediction result $\hat{y}$.
3:$M_{i}=Morgan(S),P_{i}=PubChem(S),E_{i}=PharmaERG(S)$
4:$U_{i}=[M_{i}\text{ }||\text{ }P_{i}\text{ }||\text{ }E_{i}]$
5:$FP_{i}=MLP(U_{i})$
6:Molecular Graph $G=(\mathcal{V},\mathcal{E})=GraphFunction(S)$
7:$V_{i}=InitialNode(v_{i}),E_{uv}=InitialEdge(e_{uv}),v_{i}\in\mathcal{V},e_{vu}\in\mathcal{E}$
8:for $l=1$ to $L_{1}$ do
9:     for $v_{i}\in\mathcal{V}$ do
10:         if $l==1$ then
11:              $g_{T}=GraphTransformer(V_{i})$
12:         else
13:              $g_{T}=GraphTransformer(g_{T})$
14:         end if
15:     end for
16:end for
17:for $l=1$ to $L_{2}$ do
18:     for $v_{i}\in\mathcal{V},e_{iu}\in Neightbor(v_{i})$ do
19:         if $l==1$ then
20:              $g_{A}=GraphAttention(V_{i},E_{iu})$
21:         else
22:              $g_{A}=GraphAttention(g_{A})$
23:         end if
24:     end for
25:end for
26:$G_{M}=Mixture(g_{A},g_{T})$
27:$G_{M}=MoleculeAttention(G_{M})$
28:$F_{M}=CrossAttention(FP_{i},G_{M})$
29:return $\hat{y}=\text{MLP}(F_{M})$
The general architecture of the model is shown in Figure 1. The framework integrates both molecular graph structure and fingerprint information to comprehensively capture the properties of molecules. Specifically, for the molecular graph branch, a Graph Attention Network (GAT) is employed to extract local structural information, such as function groups, by emphasizing the importance of neighboring atoms. In parallel, a modified Graph Transformer module is introduced to capture global dependencies across the entire molecular graph, enabling the model to learn long-range interactions between atoms. The output of the GAT and Graph Transformer modules are then adaptively fused through a learned weighting mechanism, allowing the model to dynamically balance local and global information. This fused graph-based representation is subsequently combined with molecular fingerprint features through a Cross-Attention layer, which facilitates deep interaction between the two modalities. The resulting joint representation is used for downstream molecular property prediction tasks.
Figure 1: The Multi-Level Fusion Graph Neural Network Architecture (A) The overview of model. (B) The architecture Mixed-information Layer.
2.3 Molecular fingerprints network
To capture comprehensive molecular-level information, we constructed a composite fingerprint representation by concatenating three widely used molecular descriptors: Morgan fingerprint 30 $M_{i}$, PubChem Fingerprint 31 $P_{i}$, and Pharmacophore ErG Fingerprint 32 $E_{i}$. Each fingerprint captures different aspects of the molecular structure and properties. Through the smiles $x_{i}$ of the molecules, we can get the corresponding molecular fingerprints. The Morgan Fingerprint encodes circular substructures, PubChem Fingerprint incorporates a variety of predefined structural keys, and ErG Fingerprint reflects pharmacophore features relevant for ligand-receptor interactions. By combining these fingerprints, we form a unified vector representation $U_{i}$ that encompasses a broad spectrum of molecular characteristics,
$U_{i}=[M_{i}\text{ }||\text{ }P_{i}\text{ }||\text{ }E_{i}].$ (1)
This concatenated fingerprint is then put into a multi-layer perceptron(MLP), which transforms it into a fixed-size embedding representing the global features of the molecule,
$FP_{i}=MLP(U_{i}).$ (2)
This embedding is subsequently used in down stream prediction tasks within out neural network framework so that it can get global information and can choose some information that is useful for predicting the molecular property.
2.4 Molecular featurization
Table 1: Initial Atomic and Bond Features
Atom Features Size Description
Atom Symbol 16 [C, N, O, F, Si, Cl, As, Se, Br, Te, I, At, others] (one-hot)
Degree 6 Number of atoms connected to the atom (one-hot)
Formal Charge 1 Electrical charge (integer)
Radical Electrons 1 Number of radical electrons (integer)
Hybridization 6 [sp, sp2, sp3, sp3d, sp3d2, other] (one-hot)
Aromaticity 1 Whether the atom is part of an aromatic system (one-hot)
Hydrogens 5 Number of hydrogens connected to the atom (one-hot)
Chirality Type 4 The chirality type of the atom (one-hot)
Ring 1 Whether the atom is in a ring (one-hot)
Ring Type 4 [3, 4, 5, 6] (one-hot)
Atomic Mass 1 Mass of the atom (integer)
Implicit Valence 7 [0, 1, 2, 3, 4, 5, 6, 7] (one-hot)
Hydrogen Acceptor 1 Whether the atom acts as a hydrogen acceptor (one-hot)
Hydrogen Donor 1 Whether the atom acts as a hydrogen donor (one-hot)
Acidic 1 Whether the atom is part of an acidic group (one-hot)
Basic 1 Whether the atom is part of a basic group (one-hot)
Bond Features Size Description
Bond Type 5 [exists, single, double, triple, aromatic] (one-hot)
Conjugation 1 Whether the bond is conjugated (one-hot)
Ring 1 Whether the bond is in a ring (one-hot)
Stereo 6 [0, 1, 2, 3, 4, 5] (one-hot)
To featurize the nodes and edges in the molecular graph, we encode them into vectors derived from a set of well-established atomic and bond characteristics. In this work, we employ 16 types of atomic features and 4 types of bond features to effectively capture the chemical identity of atoms and their local chemical environment (see Table 1). All features are encoded using one-hot vectors, except for atomic mass and formal charge, which are preserved as continuous numerical values. For atom features, since hydrogen atoms are not explicitly represented in SMILES strings, the implicit valence of atoms is used as a proxy to infer the presence of bonded hydrogens thereby preserving chemical completeness. Additionally, we incorporate whether an atom functions as a hydrogen bond donor or acceptor, and whether it is part of an acidic or basic function group, as these properties are critical for molecular interactions and biological activity. For bond features, stereo-chemical information is also included, as it reflects the spatial arrangement of atoms, which can significantly influence molecular properties and interactions. As a result, for each molecular graph $G_{i}=\{\mathcal{V}_{i},\mathcal{E}_{i}\}$, each node $v\in\mathcal{V}_{i}$ is represented by a 57-dimensional feature vector $A_{v}$, and each edge $e_{vu}\in\mathcal{E}_{i}$ is encoded with a 13-dimensional feature vector $B_{vu}$.
2.5 Molecular graph Representation
Building upon the Attentive FP framework 33, which employs graph attention networks (GAT17), our approach utilizes a similar graph attention mechanism to encode molecular graphs effectively. However, unlike Attentive FP, which leverages deep multi-layer attention to capture global dependencies across the molecule, our design intentionally employs a shallower attention stack. This enables the model to focus on local substructure patterns, which are often more informative for certain molecular properties(e.g., functional groups or local motifs). In particular, we apply only a few attention layers to preserve locality and reduce over-smoothing, allowing the model to capture neighborhood-level information more effectively. To formally describe our Graph Attention Layer, consider a molecular graph $G_{i}=\{\mathcal{V}_{i},\mathcal{E}_{i}\}$, where $\mathcal{V}_{i}$ and $\mathcal{E}_{i}$ denote the sets of atoms (nodes) and bonds(edges), respectively. The initial feature representations for atoms $A_{v}$ and bonds $B_{vu}$ are constructed using a set of domain-specific chemical descriptors, capturing both intrinsic and relational chemical properties.
Initially, each atomic feature vector $A_{V}$ is projected through a fully connected layer followed by a ReLU activation function to obtain the initial node embedding:
$h_{v}^{0}=ReLU(W_{fc1}\cdot A_{v}).$ (3)
To incorporate information from each neighboring atom $u\in N(v)$, we concatenate its atomic feature $A_{u}$ with the bond feature $B_{vu}$ yielding the neighbor representation:
$n_{u}=[A_{u}\textbf{ }||\textbf{ }B_{vu}].$ (4)
This neighbor representation is further processed through a fully connected layer with ReLU activation function
$h_{u}^{0}=ReLU(W_{fc2}\cdot n_{u}).$ (5)
After computing all initial node embeddings, we proceed to apply the $i$-th Graph Attention Layer. For each pair of connected nodes $(v,u)$, we calculate the unnormalized attention coefficient using a shared linear transformation followed by a LeakyReLU activation:
$e_{vu}^{i-1}=LeakyReLU(W_{i}[h_{v}^{i-1}||h_{u}^{i-1}]).$ (6)
These attention coefficients are then normalized over all neighbors of node $v$ using the softmax function:
$a_{vu}^{i-1}=softmax(e_{vu})=\frac{exp(e_{vu}^{i-1})}{\sum_{u\in N(v)}exp(e_{vu}^{i-1})}.$ (7)
Here, $N(v)$ denotes the set of neighboring nodes of atom $v$. Using the computed attention weights, the aggregated neighborhood representation is given by:
$c_{v}^{i-1}=elu(\sum_{u\in N(v)}a_{vu}^{i-1}\cdot W\cdot h_{u}^{i-1}).$ (8)
To enhance the expressive capacity of the model, we update the node embedding using a GRU34, which integrates the aggregated neighborhood information with the previous hidden state:
$h_{v}^{i}=GRU(C_{v}^{i-1},h_{v}^{i-1}).$ (9)
where GRU denotes the gated recurrent unit function. In our framework, the GRU acts not only as an update mechanism but also plays a critical role in refining the local representations obtained from the attention layer. While the attention mechanism enables each node to selectively aggregate information from its neighbors, the GRU further enhances this representation by recurrently integrating the newly aggregated neighborhood context with the node’s prior hidden state. This temporal integration allows the model to effectively retain and propagate locally relevant information across multiple layers, which is particularly beneficial for capturing subtle but chemically meaningful patterns in the local molecular environment. Consequently, the inclusion of the GRU improves the expressiveness and robustness of node representations, thereby enhancing performance in downstream tasks such as molecular property prediction.
While GAT captures local motifs, long-range dependencies require a global mechanism. We adopt a Transformer-based architecture enhanced with structural priors and the detailed structure of the new Transformer architecture is illustrated in Figure 2. Specifically, we integrate the graph’s normalized adjacency matrix $A$ into the self-attention weights:
$\mathcal{A}_{i}=\left(\lambda_{a}\cdot softmax\left(\frac{Q_{i}K_{i}^{T}}{\sqrt{d_{k}}}\right)+\lambda_{b}\cdot A\right)V_{i},$ (10)
where $Q_{i},K_{i},V_{i}$ are query, key, and value matrices; $\lambda_{a},\lambda_{b}$ are learnable scalars controlling the balance between learned attention and explicit graph structure.
Additionally, we replace the conventional Layer Normalization used in Transformer blocks with a novel Dynamic Tanh(DyT) 35 layer. DyT is a recently proposed alternative to traditional normalization layers, introduced as a lightweight and effective element-wise nonlinearity. Unlike Layer Normalization, which requires computing per-token statistics to rescale activations, DyT operates via a simple learnable squashing function defined as:
$DyT(x)=\gamma\cdot tanh(\alpha x)+\beta.$ (11)
where $\alpha$ is a learnable scalar that dynamically scales the input, and $\gamma,\beta$ are learnable affine parameters similar to those in normalization layers. This formulation preserves the S-shaped activation squashing effect observed in LayerNorm outputs, particularly for extreme activation values, while significantly reducing computational overhead.
After processing molecular graphs through both GAT and Transformer pathways, we obtain two complementary node-level feature streams: local representations ${H_{i}}$ from GAT layers, where each $H_{i}\in\mathbb{R}^{n\times d}$ corresponds to the $i$-th layer’s output for all $n$ atoms, and global representations $H_{T}\in\mathbb{R}^{n\times d}$ from the Transformer, encoding long-range dependencies across the molecular graph.
To integrate these two perspectives, we introduce a Mixed Information Layer that adaptively fuses the local and global features. First, we perform inter-layer aggregation over the outputs of all GAT layers. This is motivated by the observation that different GAT layers capture different neighborhood sizes (receptive fields), and aggregating them provides a more expressive local representation. The average over all GAT layers is computed as:
$Mean(H)=\frac{1}{n}\sum_{i=1}^{n}H_{i}$ (12)
where $n$ is the number of GAT layers and $H_{i}\in\mathbb{R}^{n\times d}$ is the node embedding from the $i$-th GAT layer. This produces a unified node-level representation reflecting multi-hop local information. This is then passed through a linear layer and a GeLU activation to yield the processed local feature stream:
$F_{A}=GeLU(Linear(Mean(H)))$ (13)
Here, $F_{A}\in\mathbb{R}^{n\times d}$ is the transformed local representation. The Linear layer enables feature transformation, while GeLU introduces smooth non-linearity for better generalization. In parallel, the output from the Transformer encoder $H_{T}$ is also non-linearly activated:
$F_{T}=GeLU(H_{T})$ (14)
This step ensures consistency in activation treatment across both local and global streams. $F_{T}\in\mathbb{R}^{n\times d}$ contains node-level global features learned via content-aware and adjacency-guided attention mechanisms. We combine $F_{A}$ and $F_{T}$ using a learnable scalar gate $\alpha\in[0,1]$, trained end-to-end with the model. The fusion is defined as:
$F_{Mix}=\alpha\cdot F_{A}+(1-\alpha)\cdot F_{T}$ (15)
This equation ensures an adaptive balance between local substructure sensitivity ($F_{A}$) and global topological awareness ($F_{T}$), allowing the model to modulate emphasis depending on the molecular property being predicted. For instance, properties like reactivity may benefit more from local patterns, whereas solubility may require more global context. To convert node-level features into a single molecular representation, we introduce a Virtual Super Node $H_{v}$, which serves as a global summarization anchor. It is initialized as the sum of all atomic embeddings from the fused representation:
$H_{v}=\sum_{i=1}^{n}H_{i}$ (16)
Here, $H_{i}\in\mathbb{R}^{d}$ is the embedding of the $i$-th atom in $F_{Mix}$, and $H_{v}\in\mathbb{R}^{d}$ acts as a global placeholder for the molecular context. We then compute atom-to-supernode attention scores to allow the model to weigh atoms differently, based on their relevance to the final molecular property. For each atom $i$, we compute:
$e_{vi}=LeakyReLU(W_{1}[H_{v}||H_{i}])$ (17)
where $[H_{v}||H_{i}]$ denotes concatenation of the global super node and atomic embedding, $W_{1}\in\mathbb{R}^{d\times 2d}$ is a learnable weight matrix, and $e_{vi}\in\mathbb{R}$ is an unnormalized scalar attention score reflecting the interaction between the $i$-th atom and the global context. These scores are normalized via softmax to obtain attention weights:
$a_{vi}=\frac{\exp(e_{vi})}{\sum_{j=1}^{n}\exp(e_{vj})}$ (18)
Using the attention weights $a_{vi}$, we aggregate the atom embeddings into a context vector $C_{v}$:
$C_{v}=elu\left(\sum_{i=1}^{n}a_{vi}\cdot W_{2}\cdot H_{i}\right)$ (19)
where $W_{2}\in\mathbb{R}^{d\times d}$ is a learnable transformation matrix, and $elu$ provides smooth activation to stabilize training and enhance expressiveness. $C_{v}\in\mathbb{R}^{d}$ captures a weighted summary of atomic features, focused on the most informative atoms. Finally, we update the virtual node $H_{v}$ using a GRU to integrate this attention-refined context:
$F_{v}=GRU(C_{v},H_{v})$ (20)
The Virtual Super Node $F_{v}$ serves as a global, task-adaptive representation of the entire molecule. It is subsequently passed into a MLP for downstream prediction tasks. Importantly, the introduction of this virtual node is not merely for summarization, but plays a crucial role in further refining the fused local-global information obtained from Adaptive Mixture Layer. By attending to most relevant atomic embeddings via an attention mechanism and incorporating this context through a GRU-based update, the virtual node effectively filters and distills the most predictive aspects of the mixed features. This process enhances the model’s capacity to generate expressive and discriminative molecular representations, ultimately improving its performance on various prediction tasks
Figure 2: The details of Graph Transformer Layer.
2.6 Final Representation Fusion and Output Layer
To obtain the final molecular representation for prediction, we design a fusion mechanism that effectively integrates both structural and fingerprint-based features. Specifically, the molecular graph is processed through multiple layers of extraction and filtering, capturing both global and local structural information. This process yields a comprehensive structural representation of the molecule.
In parallel, the molecular fingerprint, defined as a predefined feature vector that encodes molecular substructures, is transformed through MLP to generate a learned fingerprint representation. To enable rich interaction between these two modalities, we employ a cross-attention mechanism that jointly considers the structural and fingerprint-based representations. This module allows the model to adaptive focus on the most informative components from each source, effectively enhancing the representational capacity.
The fused output from the cross-attention module, containing integrated and refined information from both the molecular graph and fingerprint domains, is then passed through a final MLP. This layer performs the ultimate transformation required for the target prediction task producing the final prediction of the molecular property.
3 Results and Discussion
3.1 Benchmark datasets
In this study, we evaluated the performance of our model on 10 publicly available data sets from Wu et al36. These include five classification datasets: BACE, BBBP, Tox21, SIDER, and ClinTox, as well as five regression datasets including ESOL, FreeSolv, PDBbind-C, PDBbind-R and Lipophilicity. The classification tasks primarily focus on the prediction of physiology or biophysics, with the evaluation based on the ROC-AUC metric. The regression tasks target physicochemical properties and are evaluated using RMSE. These datasets span diverse molecular properties and cover a broad spectrum of molecular structures, providing a comprehensive testbed for assessing the generalization capability of our model. More information about these datasets can be found in Table 2.
Category Dataset Molecules Tasks Splitting Task type Metric Description
Biophysics PDBbind-C 168 1 Random Regression RMSE Binding affinity
PDBbind-R 3040 1 Random Regression RMSE Binding affinity
BACE 1513 1 Scaffold Classification ROC-AUC Inhibitory activity
Physiology BBBP 2053 1 Scaffold Classification ROC-AUC BBB permeability
SIDER 1427 27 Random Classification ROC-AUC Drug-related side effect
Tox21 7831 12 Random Classification ROC-AUC Compound toxicity
ClinTox 1478 2 Random Classification ROC-AUC Clinical toxicity
Physical chemistry ESOL 1128 1 Random Regression RMSE Aqueous solubility
FreeSolv 642 1 Random Regression RMSE Hydration Free energy
Lipophilicity 4200 1 Random Regression RMSE Compound LogD values
Table 2: An overview of molecular property prediction benchmark dataset.
3.2 Baselines models
To evaluate the performance of our model, we considered two groups of baseline models. The first group consists of supervised learning methods, including MoleculeNet36, D-MPNN37, ML-MPNN38, FP-GNN23, MVGNN39, LineEvo40, HimGNN41. The second group is based on self-supervised learning approaches, including GROVER6, MGSSL25, HiMOL42. A brief introduction to the aforementioned baseline model is provided below.
  • MoleculeNet36 employs a diverse set of benchmark models for molecular property prediction, encompassing traditional machine learning algorithms such as Random Forest, as well as Graph-based approaches including GCN.
  • D-MPNN37 utilizes a directed edge-based message passing framework to more effectively capture information about chemical bonds, which enhances the performance of molecular property prediction.
  • ML-MPNN38 enhances molecular property prediction by integrating sequence-level information in addition to conventional graph-level features
  • FP-GNN23 aggregates molecular fingerprint features and graph-based representations to facilitate molecular property prediction.
  • MVGNN39 introduces a novel Cross-Dependent Message Passing Scheme that facilitates mutual interaction between node and edge features, thereby improving the model’s expressive power.
  • LineEvo40 proposes a new architectural component, the LineEvo Layer, which captures hierarchical granular features from molecular structures to enhance prediction performance.
  • HimGNN41 introduces motif-level representations to extract structural information from molecular graphs, which are subsequently integrated with traditional graph features to improve the accuracy of molecular property prediction.
  • GROVER6 proposes a novel Transformer-based architecture that jointly leverages context prediction and motif prediction tasks to pretrain on large-scale molecular graphs, thereby learning transferable representations for molecular property prediction.
  • MGSSL25 enhances molecular property prediction by combining motif-level generative tasks with multi-level self-supervised pretraining on molecular graphs
  • HiMOL42 integrates node-level, motif-level, and graph-level representations, leveraging a pretraining framework to enhance performance on molecular predictions tasks.
The performance metrics of the baseline models were obtained from existing literature. Specifically, for most datasets, we adopted the baseline results reported in the HimGNN41, while for datasets under the PDBbind category, the results were referenced from the FP-GNN23. To ensure a fair comparison with these baseline models, we aligned certain experimental settings with those used in the corresponding studies.
3.3 Implementation details
We implemented our model using the RDKit and Pytorch, employing the Adam 43 optimizer to update model parameters. Hyperparameter tuning was conducted using Bayesian Optimization 44 to efficiently explore the search space. Specifically, we optimized seven key hyperparameters: (1) the number of Transformer layers $N$, (2) the number of attention heads $h$, (3) the dimensionality of each attention head $d_{k}$, (4) the output dimension of the GAT module $G_{out}$, (5) the dropout rate in the feedforward network $F_{d}$, (6) the dropout rate in the GAT module $G_{d}$, and (7) the dropout rate in the Transformer module $A_{d}$. The results of hyperparameter search are summarized in Table 3. For each hyperparameter configuration, the model was trained using 5 different random seeds, and the configuration with the highest average performance was selected as the final setting. Dataset splitting was performed according to the protocol described in Table 2. In the case of random splits, the dataset was divided into training, validation, and test sets with a ratio of 8:1:1. The validation set was used to select the final model for the better generalization ability. For each dataset, we trained our model using 10 different random seeds. All models were trained on NVIDIA GeForce RTX 4090.
Iter $N$ $h$ $d_{k}$ $G_{out}$ $G_{d}$ $F_{d}$ $A_{d}$ RMSE
1 5 13 104 140 0.50 0.55 0.25 0.639 ± 0.018
2 4 13 120 140 0.05 0.20 0.20 0.638 ± 0.017
3 3 20 120 140 0.40 0.20 0.20 0.643 ± 0.030
4 2 17 64 150 0.25 0.35 0.00 0.636 ± 0.015
5 3 19 128 140 0.35 0.55 0.20 0.655 ± 0.018
6 3 17 120 120 0.60 0.30 0.25 0.638 ± 0.024
7 3 16 64 120 0.05 0.45 0.15 0.638 ± 0.019
8 5 14 104 130 0.60 0.10 0.35 0.637 ± 0.032
9 6 19 128 160 0.15 0.35 0.15 0.643 ± 0.013
10 5 18 96 150 0.25 0.00 0.55 0.658 ± 0.031
11 4 19 120 150 0.25 0.10 0.25 0.625 ± 0.024
12 5 19 80 120 0.55 0.10 0.55 0.647 ± 0.018
13 5 14 112 150 0.35 0.10 0.30 0.653 ± 0.030
14 3 12 96 130 0.35 0.15 0.05 0.639 ± 0.018
15 4 17 96 100 0.10 0.20 0.60 0.634 ± 0.024
16 3 16 112 110 0.15 0.35 0.05 0.639 ± 0.023
17 3 19 96 110 0.50 0.05 0.50 0.610 ± 0.016
18 2 14 64 100 0.25 0.15 0.60 0.629 ± 0.018
19 2 15 64 100 0.05 0.40 0.60 0.618 ± 0.017
20 5 16 64 100 0.25 0.45 0.50 0.631 ± 0.030
Table 3: Bayesian optimization for Lipophilicity prediction task. The gray-highlighted row indicates the best hyperparameter setting.
3.4 Performance analysis
To evaluate the effectiveness of our proposed model, we conduct a comprehensive comparison several strong baselines across a suite of benchmark datasets covering both classification and regression tasks. Table 4 presents the mean and standard deviation of ROC-AUC scores on the classification benchmarks. The proposed model achieves the highest performance on three out of five datasets and ranks second on the remaining two, demonstrating its robustness and consistency across diverse molecular property prediction tasks. Similarly, Table 5 and Table 6 summarize the RMSE results for regression tasks. Our model consistently outperforms all baselines, highlighting its superior ability to capture intrinsic molecular features and the complex relationships among atoms and substructures. In summary, the experimental results on both classification and regression tasks demonstrate the strong generalization ability and effectiveness of our model. By leveraging multi-level structural information and a hierarchical representation learning framework, the proposed model is capable of capturing complex molecular features and dependencies, leading to consistently superior performance across a variety of molecular property prediction benchmarks. These findings validate the design of our approach and its potential for broad applicability in computational chemistry and drug discovery tasks.
Methods BACE BBBP SIDER Tox21 ClinTox
MoleculeNet 0.802 ± 0.038 0.877 ± 0.036 0.593 ± 0.035 0.772 ± 0.041 0.855 ± 0.037
ML-MPNN 0.819 ± 0.036 0.891 ± 0.013 0.609 ± 0.017 0.796 ± 0.021 0.865 ± 0.027
D-MPNN 0.823 ± 0.038 0.911 ± 0.048 0.610 ± 0.027 0.808 ± 0.023 0.879 ± 0.040
MVGNN 0.829 ± 0.034 0.914 ± 0.039 0.614 ± 0.029 0.811 ± 0.026 0.884 ± 0.053
FP-GNN 0.845 ± 0.028 0.910 ± 0.027 0.598 ± 0.014 0.803 ± 0.024 0.765 ± 0.038
GROVER 0.835 ± 0.044 0.911 ± 0.008 0.621 ± 0.006 0.806 ± 0.017 0.882 ± 0.013
MGSSL 0.805 ± 0.072 0.845 ± 0.011 0.591 ± 0.016 0.788 ± 0.022 0.877 ± 0.021
LineEvo 0.825 ± 0.034 0.905 ± 0.039 0.608 ± 0.011 0.835 ± 0.015 0.890 ± 0.036
HiMol 0.840 ± 0.072 0.865 ± 0.064 0.593 ± 0.022 0.815 ± 0.010 0.702 ± 0.092
HimGNN 0.856 ± 0.034 0.928 ± 0.027 0.642 ± 0.023 0.807 ± 0.017 0.917 ± 0.030
MLFGNN 0.900 ± 0.027 0.934 ± 0.014 0.647 ± 0.016 0.818 ± 0.007 0.892 ± 0.033
Table 4: Performance on classification tasks in terms of ROC-AUC. Higher is better and the best results in bold.
Methods PDBbind-C PDBbind-R
D-MPNN 1.910 ± 0.000 1.338 ± 0.000
FP-GNN 1.876 ± 0.000 1.349 ± 0.000
MLFGNN 1.816 ± 0.074 1.330 ± 0.022
Table 5: Performance on PDBbind datasets in terms of RMSE. Lower is better and the best results in bold.
Methods ESOL FreeSolv Lipophilicity
MoleculeNet 1.068 ± 0.057 2.398 ± 0.259 0.712 ± 0.041
ML-MPNN 0.905 ± 0.225 2.264 ± 0.112 0.649 ± 0.046
D-MPNN 0.972 ± 0.097 2.170 ± 0.536 0.652 ± 0.051
MVGNN 0.889 ± 0.152 2.115 ± 0.239 0.610 ± 0.059
FP-GNN 1.282 ± 0.332 2.492 ± 0.649 0.679 ± 0.152
GROVER 0.921 ± 0.092 2.026 ± 0.127 0.646 ± 0.021
MGSSL 1.935 ± 0.332 2.890 ± 0.837 0.951 ± 0.192
LineEvo 0.929 ± 0.137 2.254 ± 0.418 0.624 ± 0.011
HiMol 0.986 ± 0.231 2.731 ± 0.440 0.716 ± 0.254
HimGNN 0.870 ± 0.154 1.921 ± 0.474 0.632 ± 0.016
MLFGNN 0.641 ± 0.055 1.409 ± 0.203 0.608 ± 0.013
Table 6: Performance on regression tasks in terms of RMSE. Lower is better. Best results in bold.
3.5 Ablation study
In order to access the contribution of each component within our model architecture, we performed a series of ablation experiments targeting the Graph Transformer, the Molecular Graph representation, and the Fingerprint-based features. The ablation study results are presented in Figure 3, Figure 4, Figure 5.
In Figure 3, we investigate the impact of different configurations by replacing the original Graph Transformer module, which integrates DyT and Adjacency Matrix with three alternative variants: (1) Graph Transformer with only LN, (2) Transformer with only DyT, and (3) Transformer with LN and Adjacency Matrix. Experimental results across different datasets demonstrate that Transformer with DyT and Adjacency Matrix configuration consistently outperforms the others. These findings indicate that our proposed Transformer architecture is more effective in capturing inter-atomic relationships compared to other Transformer-based variants.
Figure 4 illustrates the results of the ablation experiments focusing on the Fingerprint module. We replaced the original concatenation of three molecular fingerprints in our model with two alternative configurations: using only the Morgan fingerprint and using the three fingerprints adopted in FP-GNN23(MACCS, PubChem, and ErG). Additionally, we performed an ablation in which the entire Fingerprint component was removed. The consistent superiority of the Morgan, PubChem, ErG fingerprints combination across all datasets demonstrates the effectiveness of combining diverse and complementary fingerprint features. In particular, the inclusion of Morgan fingerprints provides substantial gains over the MACCS-based combination used in prior models such as FP-GNN, suggesting that Morgan captures more informative substructural details. Moreover, the performance degradation observed when removing the Fingerprint component altogether underscores its importance in supplementing graph-based representation, enabling the model to leverage complementary molecular information for more accurate predictions.
Figure 5 presents the results of ablation experiments on the Molecular Graph representation. Specifically, we evaluated two configurations: one using only the GAT module and the other using the Graph Transformer module. The results show that combining GAT and Graph Transformer consistently outperforms using either module alone across all datasets. This suggests that our model effectively leverages the strengths of both components: GAT captures local substructural information such as molecular motifs and function groups, while the Graph Transformer captures global dependencies by modeling interactions between distant atoms. The complementary nature of local and global information enables the model to learn more expressive and task-relevant molecular representations.
Figure 3: The performance of MLFGNN with different Graph Transformer architecture.
Figure 4: Ablation Study on Molecular Fingerprint Features.
Figure 5: Ablation Study on Molecular Graph Module Structure.
3.6 Visualization analysis
To gain a deeper understanding of the contributions of different components in MLFGNN, we conducted interpretability analysis using SHAP (SHapley Additive exPlanation) values. SHAP is a game-theoretic approach that explains the output of machine learning models by assigning each feature an importance value for a particular prediction. It is based on Shapley values from cooperative game theory and ensure consistency and local accuracy in feature attribution 45.
Specifically, we calculated SHAP values separately for the fingerprint component and the molecular graph component, enabling us to quantify and compare their respective contributions to the model’s predictions. As shown in Figure 6, the contribution analysis across various datasets reveals that the fingerprint and molecular graph components exhibit varying levels of importance depending on the specific prediction task. In some cases, the molecular graph component plays a more dominant role, while in others, the fingerprint component contributes more significantly. This variation indicates that the two components capture different but complementary aspects of molecular information. Overall, the results suggest that combining both representations enables the model to better generalize across diverse tasks by leveraging the strengths of each modality.
To further understand the respective contributions of the Graph Attention Layer and the Graph Transformer Layer, which correspond to local and global information respectively, we analyzed the learned weighting coefficients from the Adaptive Mixture Layer. These coefficients indicate the relative influence of each component in forming the final molecular representation.
As illustrated in Figure 7, the contributions ratios between the Graph Attention and Graph Transformer components vary across datasets. This variation indicates that local and global information provide complementary insights, enabling the model to adaptively extract features that are well-suited to requirements of downstream tasks. These finding further support the effectiveness of incorporating both components in molecular representation learning.
Figure 8 and Figure 9 illustrate the molecular structures alongside atom-level feature correlation heatmaps at different stages of MLFGNN. The visualizations highlight how atomic relationships evolve before training, after the GAT layer, and after the Mixed-information layer. The heatmaps reveal how different components of the model contribute to refinement of atomic feature representations. After passing through the GAT layer, localized structural patterns such as rings and specific functional groups, are more prominently captured, reflecting the GAT’s strength in modeling local dependencies. When the Graph Transformer is further integrated integrated via the Adaptive Mixture Layer, the resulting representations incorporate additional long-range dependencies. This enables the model to better address challenges like activity cliffs, where subtle, spatially distant atomic differences significantly affect molecular properties. Overall, these visualizations demonstrate the complementary roles of local and global information in generating expressive molecular representations.
Figure 10 and Figure 11 provide visualizations of atom-level contributions to molecular graph representations for two prediction tasks. In both cases, the model assigns higher importance to chemically meaningful substructures. For instance, in Figure 10, the highlighted atoms correspond to function groups, such as aromatic rings and nitrogen-containing motifs, which are known to influence BBB permeability. Similarly, in Figure 11, atoms within hydrophobic aromatic rings of polar functional groups receive higher attention, aligning well with known determinants of lipophilicity. These observations suggest that the model is capable of capturing domain-relevant chemcial knowledge, supporting the interpretability and reliability of its learned representations.
Figure 6: Contribution percentages of Fingerprint and Graph features across classification and regression datasets.
Figure 7: Comparison of Graph Transformer and GAT contributions across various datasets.
(a) Visualization of atomic correlation patterns in the BBBP dataset
(b) Visualization of atomic correlation patterns in the ClinTox dataset
(c) Visualization of atomic correlation patterns in the Tox21 dataset
Figure 8: Molecular structures and atom feature heatmaps before and after GAT/mixed layers in some classification datasets. From left to right, each group of images shows: (1) the molecular structure, (2) the atom feature heatmap before training, (3) the atom feature heatmap after the GAT layer, and (4) the atom feature heatmap after the Mixed-information layer
(a) Visualization of atomic correlation patterns in the ESOL dataset
(b) Visualization of atomic correlation patterns in the PDB-R dataset
(c) Visualization of atomic correlation patterns in the Lipo dataset
Figure 9: Molecular structures and atom feature heatmaps before and after GAT/mixed layers in some regression datasets. From left to right, each group of images shows: (1) the molecular structure, (2) the atom feature heatmap before training, (3) the atom feature heatmap after the GAT layer, and (4) the atom feature heatmap after the Mixed-information layer.
Figure 10: Visualization of atomic contributions to the molecular graph representation for two molecules from the BBBP dataset. The color of each atom reflects its importance in determining the final graph-level embedding, with the color scale ranging from low (blue) to high (pink). The molecule on the top has a BBB permeability label of 1 (permeable), while the one on the bottom is labeled 0 (non-permeable). This visualization highlights the regions of the molecules that contribute most significantly to the prediction of blood-brain barrier permeability.
Figure 11: Visualization of atomic contributions to the molecular graph representation for two molecules from the Lipophilicity dataset. The color of each atom reflects its relative importance in forming the final molecular embedding, with pink indicating higher contribution and blue indicating lower. The molecule on the top has a measured lipophilicity value of 3.10, indicating strong lipophilicity, while the one on the bottom has a value of –0.95, indicating strong hydrophilicity. The visualization highlights the substructures that are more influential in determining the predicted lipophilicity.
4 CONCLUSIONS
In this paper, we present MLFGNN, a Multi-Level Fusion Graph Neural Network designed to enhance molecular property prediction by combining structural and multi-modal information. Our model addresses the limitations of conventional GNNs by integrating both local and global molecular features via a hybrid GAT–Graph Transformer architecture. Additionally, molecular fingerprints are incorporated through a cross-attention mechanism, enabling adaptive feature selection and improved alignment with downstream tasks. Experimental results across various benchmark datasets validate the effectiveness and generalizability of our approach. Moreover, interpretability analysis confirms that MLFGNN can identify chemically meaningful patterns, including functional groups and long-range dependencies, that are crucial for accurate prediction. Overall, our framework demonstrates the potential of combining intra-graph and inter-modal fusion strategies for robust and interpretable molecular representation learning.
5 ASSOCIATED CONTENT
5.1 Data Availability Statement
All the data sets and the source code of MLFGNN can be found at https://github.com/lhb0189/MLFGNN.
6 AUTHOR INFORMATION
6.1 Corresponding Author
Hou-biao Li — School of Mathematical Sciences, University of Electronic Science and Technology of China, Chengdu 610054, China; https://orcid.org/$0$-0002-7268-307X; Email: lihoubiao0189@163.com.
6.2 Authors
XiaYu Liu — School of Mathematical Sciences, University of Electronic Science and Technology of China, Chengdu 610054, China.
Chao Fan — College of Management Science, Chengdu University of Technology, Chengdu 610059, China.
Yang Liu — School of Mathematical Sciences, University of Electronic Science and Technology of China, Chengdu 610054, China.
6.3 Author Contributions
H.L. contributed to the study design and supervised the model construction and writing of this paper; X.L. contributed to the model conception, model construction, model training, and writing of the paper; C.F. contributed to the model construction, model training, and writing of the paper; Y.L. contributed to the model construction and writing of the paper;